Grade 12Physical Chemistry

Commercial concentrated nitric acid is 15.6 M. To prepare 10 L of 6.0 M nitric acid from it,
(a) dilute 3.846 L of 15.6 M to 10.0 L solution
(b) add 10.0 L H​​2O into 3.846 L of 15.6 M solution
(C) add 3.846 L H​​2O into 6.154 L of 15.6 M solution
(D) none of the above

Binoy John
9 Years agoGrade 12

2 Answers

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
From constant moles concept that is:
M1V1 = M2V2
15.6*V = 10*6
we get,
V = 60/15.6 => 3.846 litre is required.
Singari Sai Ashwatha
4 Years ago
From constant moles concept that is:
M1V1 = M2V2
15.6*V = 10*6
we get,
V = 60/15.6 => 3.846 litre is requires.